Against the wind a small plane flew 858 miles in 2 hours and 10 minutes. The return trip took only 1 hour and 50 minutes. What was the speed of the wind? What was the speed of the plane in still air?
Answer by stanbon(75887)

Against the wind a small plane flew 858 miles in 2 hours and 10 minutes. The return trip took only 1 hour and 50 minutes. What was the speed of the wind? What was the speed of the plane in still air?
--------------------
Against the wind DATA::
dist = 858 miles ; time = 2 1/6 = 13/6 hrs ; rate = 858/(13/6) = 396 mph
-----
With the wind DATA::
dist = 858 miles ; time = 1 5/6 = 11/6 hrs ; rate = 858/(11/6) = 468 mph
--------
Equation:
With:: p + w = 468
Against: p - w = 396
-----
Subtract to solve for "w":
2w = 72
wind = 36 mph
---------------------
Cheers,
Stan H.
---------------------